myininaya (myininaya):

Satellite this is for you! ax^2+bx+c

OpenStudy (anonymous):

factor?

myininaya (myininaya):

yes to split the middles you find two factors of a*c whose product is a*c that add up to be b say the two factors of a*c=m*n and m+n=b just replace bx=mx+nx

myininaya (myininaya):

it was bothering me

OpenStudy (anonymous):

\[a(x-\frac{-b+\sqrt{b^2-4ac}}{2a})(x-\frac{-b-\sqrt{b^2-4ac}}{2a})\]

myininaya (myininaya):

lol eww

myininaya (myininaya):

do you just not like "spilty" or do you not know how to do it

OpenStudy (anonymous):

i really have no idea.

OpenStudy (anonymous):

completing the square always works

myininaya (myininaya):

x^2+5x+6 x^2+2x+3x+6 x(x+2)+3(x+2) (x+2)(x+3) 4x^2+8x+3 4x^2+6x+2x+3 2x(2x+3)+(2x+3) (2x+3)(2x+1)

myininaya (myininaya):

two examples

OpenStudy (anonymous):

so for the 5 i was looking for two numbers that add to 5 and whose product is 6? this is just regular factoring yes?

myininaya (myininaya):

yes what about the second one what did i do?

OpenStudy (anonymous):

lets see... produce of ac is 12

OpenStudy (anonymous):

so you split 8 as 2 + 6 ?

myininaya (myininaya):

omg my little satellite is growing up

myininaya (myininaya):

yes easy!

OpenStudy (anonymous):

yeah just took my training wheels off!

myininaya (myininaya):

but quad formula never fails well for quadratics (a does not equal)

myininaya (myininaya):

except for that one hair had who said we were limited on the ways we could approach one problem

myininaya (myininaya):

do you member it? i remember something about square root

OpenStudy (anonymous):

yes and you can complete the square too. sometime called "sophie germain"

myininaya (myininaya):

sophie whos that?

myininaya (myininaya):

is she hotter than me i wonder

OpenStudy (anonymous):

for example factor \[x^4+4\]

OpenStudy (anonymous):

no she is dead since 1831

OpenStudy (anonymous):

and pretended to be a man to study math!

myininaya (myininaya):

(x^2+2i)(x^2-2i) you guys are sexist and i guess we can factor this further

OpenStudy (anonymous):

so to factor \[x^4+4\] rewrite as \[x^4+2x^2+4-2x^2\]

OpenStudy (anonymous):

oh no not over complex. watch!

myininaya (myininaya):

hey satellite if these times were like those times would you let me do work and you publish it?

myininaya (myininaya):

ok go ahead

OpenStudy (anonymous):

yes, after you did the dishes

myininaya (myininaya):

and made you a sandwich?

OpenStudy (anonymous):

that too. and fetched a beer. ok i think i made a mistake it is \[x^4+4=x^4+4x^2+4-4x^2\]

myininaya (myininaya):

you know my cat can do math and i publish his work under my name becasue you know we don't allow them to talk

OpenStudy (anonymous):

now you have the difference of two squares it is \[(x^2+2)^2-4x^2\]

OpenStudy (anonymous):

so you factor as \[(x^2+2x+2)(x^2-2x+2)\]

myininaya (myininaya):

oh! i think this is pretty

myininaya (myininaya):

very cute

OpenStudy (anonymous):

try it with \[x^4+x^2+1\]

myininaya (myininaya):

putting me on the spot how sweet

OpenStudy (anonymous):

hint is what would you add to make that a perfect square

myininaya (myininaya):

lol ok

OpenStudy (anonymous):

...

myininaya (myininaya):

x^4+2x^2+1+x^2-2x^2 =(x^2+1)^2-x^2 =(x^2+1-x)(x^2+1+x)

OpenStudy (anonymous):

i am impressed

OpenStudy (anonymous):

works for factoring trinomials as well, although can be cumbersome

OpenStudy (anonymous):

and if the leading coefficient is not one you have to work with fractions (heaven forbid)

myininaya (myininaya):

i can do math withoust sleep i haven't slept i been watching this show about british/australian teenagers who turn into mermaids whenever they touch water

myininaya (myininaya):

lol

OpenStudy (anonymous):

i saw that you were here early in the morning east coast time. must have been up very late

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Sophie_Germain i dont know she looks pretty sexy to me

myininaya (myininaya):

she has a witch nose

myininaya (myininaya):

i want to make math spells and magic

myininaya (myininaya):

ok maybe i should go to sleep before i embarrassed myself

OpenStudy (anonymous):

good night!

myininaya (myininaya):

good morning

myininaya (myininaya):

its 10:43 am

myininaya (myininaya):

thanks for teaching me something it was nice